golban:

Sales tax is calculated as a percentage of the sales price. If sales tax is 7%, what is the sales tax on clothing that costs $150? $15.70 $10.50 $9.77 $7.00

golban:

tell me the answer quick pls

justus:

Okay

golban:

pls pls pls

justus:

150 * 0.07=

golban:

10.50 or 10.5

golban:

thx

justus:

That's right

golban:

wait

justus:

Yes?

golban:

ok

golban:

It costs $3.55 to make a sandwich at the local deli shop. To make a profit, the deli sells it at a price that is 160% of the cost. The sandwich sells for $___.

justus:

3.55 * 1.6=

justus:

I think we already did this one

golban:

o sorry

golban:

this one

golban:

oanne had $115 to spend at the mall. She spent 23% of that money on a shirt. Joanne spent $___ on the shirt.

justus:

Okay

justus:

0.23(115) =

golban:

so what is it

justus:

Calculate it in the calculator

golban:

26.45

justus:

Because I got a warning for giving you the answer I am only supposed to help you get to the answer I'm sorry

justus:

Correct :)

golban:

thx one more

justus:

Sure

golban:

James bought 12 cookies and ate 3 of them. He ate ____% of the cookies.

golban:

0.36

justus:

(3/12) = .25 x 100 =

golban:

sorry is it 25

golban:

At Southside Middle School, 20% of the students play an instrument. There are 305 students. How many students play an instrument? 20 students 61 students 244 students 285 students

justus:

Yes ma'am/ sir

golban:

in a guy

golban:

but what about this

justus:

0.20 * 305=

golban:

At Southside Middle School, 20% of the students play an instrument. There are 305 students. How many students play an instrument? 20 students 61 students 244 students 285 students

justus:

I'm sorry

justus:

0.20 * 305=

golban:

61 letter b

justus:

Yes sir

golban:

After a concert for 1,600 people, only 4 people said they would not go to see the band again. What percent of the people who went to the concert said they would not go to see the band again? 0.04% 0.25% 4% 25%

golban:

two more

golban:

promise

justus:

Of course I can help you

golban:

thx

justus:

4/1600 * 100 = 4/16 =

golban:

what do i do

justus:

Calculate 4/16

golban:

4 times 6

justus:

No it will be 4/16

justus:

Do it in the calculator

golban:

0.67

justus:

You will get the answer

justus:

No

golban:

what means 4/6

golban:

sorry 4/16

justus:

You just have to put 4/16 in the calculator and you get the answer

justus:

It is 0.25

golban:

oh thank you

justus:

No problem :)

golban:

Sales tax is calculated as a percentage of the sales price. If sales tax is 5%, what is the sales tax on clothing that costs $170? $8.50 $10.50 $15.00 $17.50

justus:

Be careful the answer is 0.25 not 25%

justus:

Okay

golban:

ok

justus:

0.05($170)=

golban:

what?

justus:

0.05 * 170=

justus:

Do it in the calculator I suppose

AngeI:

For future references questions should try to be limited to one per post Also by Justus explaining it a few times you should try to do it on your own so that you may learn better c:

justus:

Thank you angel

justus:

Yes correct

justus:

No

justus:

I'm here if you need any more help, good luck :)

justus:

No problem